The product of two consecutive odd integer is 1 less than twice their sum. Find the integers.
Answer by CRandunu(8)

let first odd integer be x and next odd integer be (x+2)
given
x(x+2) = 2{ x+(x+2)}-1
x^2+2x =4x+3
x^2-2x-3=0
(x+1)(x-3) = 0
x=-1 or x=3 but x cant be -1 because given x is an odd integer...hence x=3
then x+2= 3+2=5
hence two required odd integers are 3 and 5